OpenStudy (anonymous):

Add. Express your answer in simplest terms. 2/3+2/7

OpenStudy (anonymous):

First step: Find a common denominator.

OpenStudy (anonymous):

the LCD is 27

OpenStudy (anonymous):

No it isn't. 7 doesn't go into 27. Just multiply the denominators for this one.

OpenStudy (anonymous):

21

OpenStudy (anonymous):

Yeah, now just get each fraction so that it has a denominator of 21

OpenStudy (anonymous):

what do u mean

OpenStudy (anonymous):

Well how many times does 3 go into 21?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

So do \[\large \sf \frac{2}{3} \times \frac{7}{7}\] and that will get the denominator to 21

OpenStudy (anonymous):

14/21

OpenStudy (anonymous):

Good. Now do the same process for 2/7

OpenStudy (anonymous):

\[2/7 \times 3/3\]

OpenStudy (anonymous):

Yep, which would get you what?

OpenStudy (anonymous):

6/21

OpenStudy (anonymous):

Good. So now they have the same denominator. Now we can add them \[\large \sf \frac{14}{21}+\frac{6}{21}=?\]

OpenStudy (anonymous):

20/21

OpenStudy (anonymous):

Correct, and that would be your answer